solve the differential equation by the method of variation of parameters d²y/dx²+9y=sec3x


1
Expert's answer
2021-12-28T12:02:14-0500

The given differential equation can be written as;

"(D^2+9)y=sec3x , where\ D=\frac{d}{dx}"

Auxiliary equation is "(D^2+9)=0"

"\implies D=\pm3i"

Hence, C.F is,

"y_c=C_1\ cos3x+C_2\ sin3x"

Now we take "y_1=cos 3x, y_2=sin3x"

"w=\begin{vmatrix}
 y_1 & y_2 \\
 y_1' & y_2'
\end{vmatrix}=\begin{vmatrix}
 cos3x & sin3x \\
 -3sin3x & 3cos3x
\end{vmatrix}=3"

Particular integral is,

"y_p=-y_1\int \frac{y_2X}{w}dx+y_2\int \frac{y_1X}{w}dx"

"=-cos3x\int \frac{sin3xsec3x}{3}dx+sin3x\int \frac{cos3xsec3x}{3}dx"

"=-\frac{cos3x}{3}\int tan3xdx+\frac{sin3x}{3}\int dx"

"=\frac{cos 3x\ log(cos3x)}{9}+\frac{x\ sin3x}{3}"

Hence, the general solution is;

y(x)="y_c+y_p"

"y(x)=C_1\ cos3x+C_2\ sin3x+\frac{cos 3x\ log(cos3x)}{9}+\frac{x\ sin3x}{3}"